# Addressing the Infinite Variance Problem in Fermionic Monte Carlo Simulations: Retrospective Error Remediation and the Exact Bridge Link Method

### v1

@@ II.1 Infinite variance problem and heavy-tailed distribution
Figure 1: (a) Energy estimates from 50 independent DQMC runs. Error bars indicate twice the SEM, corresponding to a conventional $\sim$95.4confidence interval. Orange points denote cases where the true energy value (black horizontal line, obtained using the exact bridge link method described in Sec. III) lies outside the estimated confidence interval. (b) Variance of the energy estimator versus sample size for the first 5 independent runs, with each color indicating one run.
@@
Under standard statistical assumptions, the error bars correspond to an approximately 95.4% confidence interval for the sample mean. However, as Fig. 1(a) shows, in 6 out of 50 independent runs, the true energy value lies outside the reported confidence interval—almost three times the expected $\sim$4.6failure rate. It is also noteworthy that several runs exhibit significantly larger error bars compared to others — even though they all use the same number of samples — and in those cases, the sample mean deviates substantially from the true value. As illustrated in Fig. 1(b), the sample variance $s^{2}\equiv\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}$ fails to converge with increasing sample size $N$, but instead grows erratically, characteristic of the infinite variance problem.
@@
we extract a characteristic exponent $\alpha=1.51\pm 0.04$ and a tail coefficient $c=5.2\pm 1.2$, which is displayed in Fig. 2(b). A distribution with a power-law tail is commonly referred to as heavy-tailed, with the parameter $\alpha$ known as the characteristic exponent. When $21<\alpha<$ 2, the distribution has a finite mean but divergent variance, while for $0<\alpha\leq 1$ even the mean is undefined.
The present data falls into the $21<\alpha<$ 2 regime. A direct consequence of this is the breakdown of the central limit theorem (CLT), which requires finite variance as a prerequisite. As illustrated in Fig. 2(c), the distribution of the sample mean, defined as $\bar{X}_{N_{\text{bin}}}\equiv\frac{1}{N_{\text{bin}}}\sum_{i=1}^{N_{\text{bin}}}X_{i}$, does not converge to a normal distribution, even for large ${N_{\text{bin}}}$. This breakdown of the CLT is precisely why the conventional standard error estimator fails: without finite variance, the usual relationship between sample size and uncertainty no longer holds, leading to unreliable or misleading confidence intervals.
@@
where $\mathbb{P}(X>x)$ and $\mathbb{P}(X<-x)$ denote the cumulative probabilities that $X$ exceeds $x$ and is less than $-x$, respectively. Assume further that $1<\alpha<2,c^{-}\geq 0,c^{+}\geq 0$ and $c^{-}+c^{+}>0$. Then, the GCLT states that the properly rescaled sample mean converges in distribution to a stable law [38, 36, 37]:
@@ II.2 Tail-aware error estimation
If the CLT holds, this estimator converges to $\sigma/\sqrt{N}$ in the limit $N\rightarrow\infty$, where $\sigma$ is the variance of the underlying distribution. Multiplying the SEM by two yields an approximate $95.4\%$ confidence interval for the sample mean — commonly referred to as the two-sigma confidence interval. However, for heavy-tailed distributions with $21<\alpha<$ 2, these conventional expressions become invalid because the variance $\sigma^{2}$ is formally divergent. Counterintuitively, though, $\text{SE}_{N}$ does not diverge as the sample size goes to infinity. In the heavy-tailed case, one can show that the appropriately rescaled SEM converges in distribution to a nontrivial limiting form:
@@
    From the sorted $T_{N}(i)$, extract the lower and upper quantiles at levels $\epsilon/2$ and $21-\epsilon/$ 2, respectively. These quantiles estimate the scaling factors $\xi_{\downarrow},\xi_{\uparrow}$.
@@
In Fig. 3(b), we illustrate the dependence of the rescaling factors $\xi_{\uparrow}$ and $\xi_{\downarrow}$ on the characteristic exponent $\alpha$ and the number of bins $N$. As evident from the plot, both factors converge to asymptotic values that depend solely on $\alpha$ (with $\beta$ fixed) in the limit $N\to\infty$. As $\alpha$ approaches 2, these rescaling factors recover the conventional value of 2, consistent with the standard case of finite variance under the CLT. For $\alpha<2$, however, the rescaled error bars exceed the conventional ones, reflecting the underestimation of statistical uncertainty when using standard error estimation in the presence of heavy-tailed distributions. Even then, when $21<\alpha<$ 2, the distribution of the studentized sample mean converges to a limiting distribution with Gaussian-like tails [43]. This helps to explain why results in the literature from calculations such as DQMC, with infinite variance, can seem “reasonable” even with error estimates based on the “standard” approach. The true confidence intervals differ primarily by a scaling factor, which itself converges as the confidence level increases.
@@ III.1 The DQMC algorithm and the origin of the infinite variance problem
|$\ket{\psi_{G}}=\lim_{\Theta\rightarrow\infty}e^{-\frac{\Theta}{2}\hat{H}}\ket{\psi_{T}},$ (6)
where |$\ket{\psi_{T}}$ is a chosen trial wavefunction, typically taken to be a Slater determinant. The expectation value of the observable $\hat{O}$ is calculated as
@@
Ignoring the Trotter error due to finite time-step size $\Delta\tau$, the ground-state wavefunction in DQMC can be interpreted as a superposition of Slater determinants defined over the configuration space of auxiliary fields. Specifically, we can write: |$\ket{\psi_{G}}\simeq\sum_{\bm{x}_{L}}\ket{\phi(\bm{x}_{L})}$ where each configuration $\bm{x}_{L}=\{\bm{x}_{1},\cdots,\bm{x}_{L_{\tau}/2}\}$ defines a Slater determinant |$\ket{\phi(\bm{x}_{L})}=\prod_{i=1}^{L_{\tau}/2}e^{-\frac{\Delta\tau}{2}\hat{K}}e^{\hat{v}(\bm{x}_{i})}e^{-\frac{\Delta\tau}{2}\hat{K}}\ket{\psi_{0}}$. Similarly for the right side: |$\ket{\psi_{G}}\simeq\sum_{\bm{x}_{R}}\ket{\phi(\bm{x}_{R})}$, with $\bm{x}_{R}=\{\bm{x}_{L_{\tau}/2+1},\cdots,\bm{x}_{L_{\tau}}\}$. Then
$\braket{\hat{O}}=\frac{\sum_{\bm{x}_{L},\bm{x}_{R}}\bra{\phi(\bm{x}_{L})}\hat{O}\ket{\phi(\bm{x}_{R})}}{\sum_{\bm{x}_{L},\bm{x}_{R}}\braket{\phi(\bm{x}_{L})}{\phi(\bm{x}_{R})}}=\frac{\sum_{\bm{x}}w(\bm{x})O_{\text{loc}}(\bm{x})}{\sum_{\bm{x}}w(\bm{x})},$ (9)
where $w(\bm{x})\equiv\braket{\phi(\bm{x}_{L})}{\phi(\bm{x}_{R})}$ is the Boltzmann weight that can be calculated by a determinant, and a local observable $O_{\text{loc}}$ is defined as
$O_{\text{loc}}(\bm{x})\equiv\frac{\bra{\phi(\bm{x}_{L})}\hat{O}\ket{\phi(\bm{x}_{R})}}{\braket{\phi(\bm{x}_{L})}{\phi(\bm{x}_{R})}}.$ (10)
@@
where $i$ is an integer that runs through the lattice and labels the sites, $i,j\right>$ denotes a sum over nearest-neighbor sites, $t$ is the hopping amplitude (set to 1 throughout this paper), and $U$ is the onsite Hubbard interaction strength. The index $\sigma=\uparrow,\downarrow$ labels the spin. For the interaction term $U>0$, there is an exact discrete HS transformation [45]:
@@
Figure 4: Nodal structure of the spin-resolved weight and local energy distribution. Panels (a) and (b) show ${\omega}_{\uparrow}(\bm{x})$ as colormaps in two planar cuts in the space of auxiliary-field configurations. A configuration is first identified by detecting a sign change during the simulation. The last update causing the sign change occurred with the particular field $x_{i}$, which is shown as the horizontal axis in both panels ($\Delta x_{i}$ represents the displacement from the node position). In (a), the plane is $\Delta x_{i}$-$\Delta x_{j}$, where site $j$ is a near-neighbor of $i$. In (b), the plane is $\Delta x_{i}$-$\Delta x_{k}$, where $k$ is the furthest site from $i$. Each inset shows ${\omega}_{\uparrow}(\bm{x})$ plotted along the black dashed line indicated in the main plot. Panel (c) shows the distribution of local energies for four interaction strengths. Dashed lines represent power-law fits of the form $y=\frac{\alpha c}{(x_{0}-x)^{\alpha+1}}$. The resulting characteristic exponents are shown in the figure and the tail coefficients are $0.7\pm 0.4,4.1\pm 1.5,15.4\pm 3.5,(1.8\pm 0.4)\times 10^{2}$ for $U=-2,-3,-4,-8$, respectively. The system is a $6\times 6\times 2$ honeycomb lattice Hubbard model at half filling. The site $i$ is $(5,5,A)$, while $j=(5,4,B)$ and $k=(2,2,A)$ ($A$ and $B$ are sublattice indices). Each histogram in (c) is from $10^{7}$ data points. Each ${\omega}_{\uparrow}(\bm{x})$ is normalized by its maximum value within the plotted region in (a) and (b).
@@
In the numerator, we can rewrite: $w(\bm{x})\,O^{2}_{\text{loc}}(\bm{x})=\frac{\bra{\phi(\bm{x}_{L})}\hat{O}\ket{\phi(\bm{x}_{R})}^{2}}{w(\bm{x})}$ which clearly diverges as $w(\bm{x})$ approaches zero. If this divergence is sufficiently strong, the variance becomes infinite.
@@ III.2 The exact bridge link method
In the bridge link method, sampling from the modified distribution $\tilde{w}\propto\langle\phi_{L}|e^{-\delta\tau\hat{H}}|\phi_{R}\rangle$ is performed by inserting another time step in the conventional DQMC simulation, and then discarding the extra dimension, i.e., the additional auxiliary-field associated with the time slice ${}^{-\delta\tau\hat{H}}$, in the evaluation in Eq. (15). It is easy to see that $\tilde{w}$ remains sign-problem-free and removes the infinite variance problem, as it can be written as
@@
Figure 5: Illustrative results from the exact bridge link method. (a) DQMC results for the same system as in Fig. 1(a). Note that the vertical axis scale has been shrunk by more than 15×15$\times$15 × from Fig. 1(a). Identical simulation parameters are used, except that the bridge link method is employed with bridge operator parameter $\theta=1$ (Eq. (19)). Error bars are estimated using the Jackknife method. (b) Scatter plot of $|E_{\text{loc}}|$ versus $\Lambda_{\text{loc}}$. While both quantities can diverge in magnitude, their ratio converges to a constant, indicated by the dashed line. (c),(d) Probability densities of the ratio $E_{\text{loc}}/\Lambda_{\text{loc}}$ and of $\Lambda_{\text{loc}}^{-1}$, respectively.
@@
where $m$ is the number of the sign-free operators, and $\hat{v}_{i}$ are one-body operators with $O(1)$ support. Then the bridge value can be calculated in $O(m)$ time. In Eq. (20), the operators ${\hat{A}}$, ${\hat{B}}$, and ${\hat{v}_{i}}$ are all required to satisfy the same symmetries as the decomposed form of $e^{-\delta\tau\hat{H}}$, i.e., commute with $T_{1}$ and $T_{2}$, so as to maintain the sign-problem-free nature of the path. The main steps of the exact bridge link method are outlined as follows:
@@
    When reaching the bridge time slice, compute the propagated wavefunctions: |$\ket{\phi_{R}^{\prime}}=e^{\hat{B}}\ket{\phi_{R}}$, |$\ket{\phi_{L}^{\prime}}=e^{\hat{A}^{\dagger}}\ket{\phi_{L}}$ and evaluate the overlap ratio: $r=\frac{\braket{\phi_{L}^{\prime}}{\phi_{R}^{\prime}}}{\braket{\phi_{L}}{\phi_{R}}}$.
@@
    Construct the Green’s function from |$\ket{\phi_{L,R}^{\prime}}$ , and use the matrix determinant lemma to compute the local bridge weights: $w_{\text{loc}}(i)=\frac{\braket{\phi_{L}^{\prime}}{e^{\hat{v}_{i}}}{\phi_{R}^{\prime}}}{\braket{\phi_{L}^{\prime}}{\phi_{R}^{\prime}}}$.
@@
Figure 6: Comparison between the exact and original bridge link methods. (a) Energy obtained versus the size of the bridge link, $\delta\tau$. (b) Zoom-in on the energy differences between sub-MC results, the exact bridge link result, and the overall sample mean ($E_{0}$) obtained from the exact bridge method. Data points where $E_{0}$ lies outside the error bar are shown with reduced transparency for emphasis. (c) Finite-sample standard error of the mean, $\text{SE}_{N}$, for different bridge methods and time steps. The system is the attractive Hubbard model on a $12\times 12$ lattice with 144 electrons and interaction strength $U=-8$. The simulation parameters are $\Theta=20,\Delta\tau=0.05$. Each data point represents $10^{5}$ Monte Carlo sweeps, with the first quarter discarded to ensure thermalization. In the sub-MC calculations for the original bridge link method, $1000$ samples are used.
@@
$\hat{\Lambda}^{\prime}(\delta\tau)\propto\frac{1}{N_{\text{site}}}\sum_{i}\left(\sum_{x=\pm 1}e^{-\frac{\delta\tau}{2}\hat{K}}e^{\mathrm{i}\theta x(\hat{n}_{i\uparrow}-\hat{n}_{i\downarrow})}e^{-\frac{\delta\tau}{2}\hat{K}}\right),$ (21)
where $\theta$ is chosen as $\theta=\arccos\left(\frac{4+N_{\text{site}}U\delta\tau}{4-N_{\text{site}}U\delta\tau}\right)$, to approximate the evolution operator $e^{-\delta\tau\hat{H}}$ used in the original bridge link method. In Fig. 6, we compare the exact and original bridge link methods. The problem studied is the attractive Hubbard model on a larger $12\times 12$ lattice with a stronger interaction strength of $U=-8$, providing a stringent test case with a pronounced infinite variance problem. By varying the bridge time step $\delta\tau$, we observe that the original method—based on a second-order expansion—exhibits significant deviation as $\delta\tau$ increases. To improve accuracy, we also employ the sub-MC approach with 1000 samples to estimate the bridge value in the original method. While this greatly improves the results, a close-up view as in Fig. 6(b) still reveals a systematic bias that grows with increasing $\delta\tau$. In Fig. 6(c), we plot the SEM as a function of $\delta\tau$. For the original method, the SEM initially decreases with increasing $\delta\tau$ but then increases again. In contrast, the exact bridge method exhibits much more stable behavior across a wide range of time steps.

### v2

@@ II.1 Infinite variance problem and heavy-tailed distribution
Figure 1: (a) Energy estimates from 50 independent DQMC runs. Error bars indicate twice the SEM, corresponding to a conventional $\sim$95.4% confidence interval. Orange points denote cases where the true energy value (black horizontal line, obtained using the exact bridge link method described in Sec. III) lies outside the estimated confidence interval. (b) Variance of the energy estimator versus sample size for the first 5 independent runs, with each color indicating one run.
@@
Under standard statistical assumptions, the error bars correspond to an approximately 95.4% confidence interval for the sample mean. However, as Fig. 1(a) shows, in 6 out of 50 independent runs, the true energy value lies outside the reported confidence interval—almost three times the expected $\sim$4.6% failure rate. It is also noteworthy that several runs exhibit significantly larger error bars compared to others — even though they all use the same number of samples — and in those cases, the sample mean deviates substantially from the true value. As illustrated in Fig. 1(b), the sample variance $s^{2}\equiv\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}$ fails to converge with increasing sample size $N$, but instead grows erratically, characteristic of the infinite variance problem.
@@
we extract a characteristic exponent $\alpha=1.51\pm 0.04$ and a tail coefficient $c=5.2\pm 1.2$, which is displayed in Fig. 2(b). A distribution with a power-law tail is commonly referred to as heavy-tailed, with the parameter $\alpha$ known as the characteristic exponent. When $1<\alpha<2$, the distribution has a finite mean but divergent variance, while for $0<\alpha\leq 1$ even the mean is undefined.
The present data falls into the $1<\alpha<2$ regime. A direct consequence of this is the breakdown of the central limit theorem (CLT), which requires finite variance as a prerequisite. As illustrated in Fig. 2(c), the distribution of the sample mean, defined as $\bar{X}_{N_{\text{bin}}}\equiv\frac{1}{N_{\text{bin}}}\sum_{i=1}^{N_{\text{bin}}}X_{i}$, does not converge to a normal distribution, even for large $N_{\text{bin}}$}. This breakdown of the CLT is precisely why the conventional standard error estimator fails: without finite variance, the usual relationship between sample size and uncertainty no longer holds, leading to unreliable or misleading confidence intervals.
@@
where $\mathbb{P}(X>x)$ and $\mathbb{P}(X<-x)$ denote the cumulative probabilities that $X$ exceeds $x$ and is less than x-x, respectively. Assume further that $1<\alpha<2,c^{-}\geq 0,c^{+}\geq 0$ and $c^{-}+c^{+}>0$. Then, the GCLT states that the properly rescaled sample mean converges in distribution to a stable law [38, 36, 37]:
@@ II.2 Tail-aware error estimation
If the CLT holds, this estimator converges to $\sigma/\sqrt{N}$ in the limit $N\rightarrow\infty$, where $\sigma$ is the variance of the underlying distribution. Multiplying the SEM by two yields an approximate $95.4\%$ confidence interval for the sample mean — commonly referred to as the two-sigma confidence interval. However, for heavy-tailed distributions with $1<\alpha<2$, these conventional expressions become invalid because the variance $\sigma^{2}$ is formally divergent. Counterintuitively, though, $\text{SE}_{N}$ does not diverge as the sample size goes to infinity. In the heavy-tailed case, one can show that the appropriately rescaled SEM converges in distribution to a nontrivial limiting form:
@@
    From the sorted $T_{N}(i)$, extract the lower and upper quantiles at levels $\epsilon/2$ and $1-\epsilon/2$, respectively. These quantiles estimate the scaling factors $\xi_{\downarrow},\xi_{\uparrow}$.
@@
In Fig. 3(b), we illustrate the dependence of the rescaling factors $\xi_{\uparrow}$ and $\xi_{\downarrow}$ on the characteristic exponent $\alpha$ and the number of bins $N$. As evident from the plot, both factors converge to asymptotic values that depend solely on $\alpha$ (with $\beta$ fixed) in the limit $N\to\infty$. As $\alpha$ approaches 2, these rescaling factors recover the conventional value of 2, consistent with the standard case of finite variance under the CLT. For $\alpha<2$, however, the rescaled error bars exceed the conventional ones, reflecting the underestimation of statistical uncertainty when using standard error estimation in the presence of heavy-tailed distributions. Even then, when $1<\alpha<2$, the distribution of the studentized sample mean converges to a limiting distribution with Gaussian-like tails [43]. This helps to explain why results in the literature from calculations such as DQMC, with infinite variance, can seem “reasonable” even with error estimates based on the “standard” approach. The true confidence intervals differ primarily by a scaling factor, which itself converges as the confidence level increases.
@@ III.1 The DQMC algorithm and the origin of the infinite variance problem
[…]
$\braket{\hat{O}}=\frac{\braket{\psi_{T}|e^{-\frac{\Theta}{2}\hat{H}}\hat{O}e^{-\frac{\Theta}{2}\hat{H}}|\psi_{T}}}{\braket{\psi_{T}|e^{-\Theta\hat{H}}|\psi_{T}}}.$ (7)
@@
[…]
where $w(\bm{x})\equiv\braket{\phi(\bm{x}_{L})|\phi(\bm{x}_{R})}$ is the Boltzmann weight that can be calculated by a determinant, and a local observable $O_{\text{loc}}$ is defined as
[…]
@@
where $i$ is an integer that runs through the lattice and labels the sites, $\left<i,j\right$> denotes a sum over nearest-neighbor sites, $t$ is the hopping amplitude (set to 1 throughout this paper), and $U$ is the onsite Hubbard interaction strength. The index $\sigma=\uparrow,\downarrow$ labels the spin. For the interaction term $U>0$, there is an exact discrete HS transformation [45]:
@@
Figure 4: Nodal structure of the spin-resolved weight and local energy distribution. Panels (a) and (b) show ${\omega}_{\uparrow}(\bm{x})$ as colormaps in two planar cuts in the space of auxiliary-field configurations. A configuration is first identified by detecting a sign change during the simulation. The last update causing the sign change occurred with the particular field $x_{i}$, which is shown as the horizontal axis in both panels ($\Delta x_{i}$ represents the displacement from the node position). In (a), the plane is $\Delta x_{i}$-$\Delta x_{j}$, where site $j$ is a near-neighbor of $i$. In (b), the plane is $\Delta x_{i}$-$\Delta x_{k}$, where $k$ is the furthest site from $i$. Each inset shows ${\omega}_{\uparrow}(\bm{x})$ plotted along the black dashed line indicated in the main plot. Panel (c) shows the distribution of local energies for four interaction strengths. Dashed lines represent power-law fits of the form $y=\frac{\alpha c}{(x_{0}-x)^{\alpha+1}}$. The resulting characteristic exponents are shown in the figure and the tail coefficients are $0.7\pm 0.4,4.1\pm 1.5,15.4\pm 3.5,(1.8\pm 0.4)\times 10^{2}$ for U=2,3,4,8U=-2,-3,-4,-8, respectively. The system is a $6\times 6\times 2$ honeycomb lattice Hubbard model at half filling. The site $i$ is $(5,5,A)$, while $j=(5,4,B)$ and $k=(2,2,A)$ (AA and $B$ are sublattice indices). Each histogram in (c) is from $10^{7}$ data points. Each ${\omega}_{\uparrow}(\bm{x})$ is normalized by its maximum value within the plotted region in (a) and (b).
@@
[…]
@@ III.2 The exact bridge link method
In the bridge link method, sampling from the modified distribution $\tilde{w}\propto\langle\phi_{L}|e^{-\delta\tau\hat{H}}|\phi_{R}\rangle$ is performed by inserting another time step in the conventional DQMC simulation, and then discarding the extra dimension, i.e., the additional auxiliary-field associated with the time slice $^{}^{-\delta\tau\hat{H}}$, in the evaluation in Eq. (15). It is easy to see that $\tilde{w}$ remains sign-problem-free and removes the infinite variance problem, as it can be written as
@@
Figure 5: Illustrative results from the exact bridge link method. (a) DQMC results for the same system as in Fig. 1(a). Note that the vertical axis scale has been shrunk by more than $15\times$ from Fig. 1(a). Identical simulation parameters are used, except that the bridge link method is employed with bridge operator parameter $\theta=1$ (Eq. (19)). Error bars are estimated using the Jackknife method. (b) Scatter plot of $|E_{\text{loc}}|$ versus $\Lambda_{\text{loc}}$. While both quantities can diverge in magnitude, their ratio converges to a constant, indicated by the dashed line. (c),(d) Probability densities of the ratio $E_{\text{loc}}/\Lambda_{\text{loc}}$ and of $\Lambda_{\text{loc}}^{-1}$, respectively.
@@
where $m$ is the number of the sign-free operators, and $\hat{v}_{i}$ are one-body operators with $O(1)$ support. Then the bridge value can be calculated in $O(m)$ time. In Eq. (20), the operators $^{\hat{A}}$, $^{\hat{B}}$, and $\hat{v}_{i}$} are all required to satisfy the same symmetries as the decomposed form of $e^{-\delta\tau\hat{H}}$, i.e., commute with $T_{1}$ and $T_{2}$, so as to maintain the sign-problem-free nature of the path. The main steps of the exact bridge link method are outlined as follows:
@@
[…]
@@
[…]
@@
Figure 6: Comparison between the exact and original bridge link methods. (a) Energy obtained versus the size of the bridge link, $\delta\tau$. (b) Zoom-in on the energy differences between sub-MC results, the exact bridge link result, and the overall sample mean ($E_{0}$) obtained from the exact bridge method. Data points where $E_{0}$ lies outside the error bar are shown with reduced transparency for emphasis. (c) Finite-sample standard error of the mean, $\text{SE}_{N}$, for different bridge methods and time steps. The system is the attractive Hubbard model on a $12\times 12$ lattice with 144 electrons and interaction strength U=8U=-8. The simulation parameters are $\Theta=20,\Delta\tau=0.05$. Each data point represents $10^{5}$ Monte Carlo sweeps, with the first quarter discarded to ensure thermalization. In the sub-MC calculations for the original bridge link method, 10001000 samples are used.
@@
[…]
where $\theta$ is chosen as $\theta=\arccos\left(\frac{4+N_{\text{site}}U\delta\tau}{4-N_{\text{site}}U\delta\tau}\right)$, to approximate the evolution operator $e^{-\delta\tau\hat{H}}$ used in the original bridge link method. In Fig. 6, we compare the exact and original bridge link methods. The problem studied is the attractive Hubbard model on a larger $12\times 12$ lattice with a stronger interaction strength of U=8U=-8, providing a stringent test case with a pronounced infinite variance problem. By varying the bridge time step $\delta\tau$, we observe that the original method—based on a second-order expansion—exhibits significant deviation as $\delta\tau$ increases. To improve accuracy, we also employ the sub-MC approach with 1000 samples to estimate the bridge value in the original method. While this greatly improves the results, a close-up view as in Fig. 6(b) still reveals a systematic bias that grows with increasing $\delta\tau$. In Fig. 6(c), we plot the SEM as a function of $\delta\tau$. For the original method, the SEM initially decreases with increasing $\delta\tau$ but then increases again. In contrast, the exact bridge method exhibits much more stable behavior across a wide range of time steps.